the kings won two less than three times as many games as they lost. they played 82 games.. how many wins and losses did the team have?
Answer by josmiceli(19441)

W + L = 82
W = 3L -2
substitute 3L - 2 for W in the first equation
3L -2 + L = 82
4L = 84
L = 21
W = 3 * 21 - 2
W = 63 - 2
W = 61
CHECK:
61 + 21 = 82
TRUE
61 = 3(21) - 2
61 = 63 - 2
TRUE
61